​Find the equation to the locus of a point which moves such that it’s distance from A(3,5) is equal to ‘X’ Axis.

Grade:11

1 Answers

siddharth gupta
9 years ago
The language of the question is not so clear.A possibility is that distance from (3,5) is equal to distance from x-axis.
accordingly:(h-3)2+(k-5)2=k2
on solving h2-6h-10k+34=0
or x2-6x-10y+34=0
with regards
siddharth gupta